OpenStudy (anonymous):

f(x) =1/(1-x)^2 obtain f ' (x) using the defining formula

OpenStudy (kirbykirby):

You could use the quotient rule, but i think the chain rule is easier... convert the fraction to a negative power and use the chain rule. \[ \frac{1}{(1-x)^2}=(1-x)^{-2}\] Chain rule: \[\frac{d}{dx}f(x)= -2(1-x)^{-3}(-1)=2(1-x)^{-2}=\frac{2}{(1-x)^2}\]

OpenStudy (anonymous):

no its the one with (f(x+h)-f(x))/h

OpenStudy (kirbykirby):

oops i changed the 3 into a 2 half-way through

OpenStudy (kirbykirby):

ooh the long way

OpenStudy (anonymous):

Lol

OpenStudy (kirbykirby):

\[\lim_{h\rightarrow 0} \frac{f(x+h)-f(x)}{h}\\ =\lim_{h\rightarrow 0}\large\frac{\frac{1}{(1-(x+h))^2}-\frac{1}{(1-x)^2}}{h}\\ \, \\ =\lim_{h\rightarrow 0}\large \frac{\frac{1}{((1-x)-h)^2}-\frac{1}{(1-x)^2}}{h}\\ \, \\ =\lim_{h\rightarrow 0}\large \frac{\frac{(1-x)^2}{(1-x)^2((1-x)-h)^2}-\frac{((1-x)-h)^2}{(1-x)^2((1-x)-h)^2}}{h}\\ \, \\ =\lim_{h\rightarrow 0} \large \frac{\frac{(1-x)^2-((1-x)-h)^2}{(1-x)^2((1-x)-h)^2}}{h}\\ \, \\ =\lim_{h\rightarrow 0}\frac{(1-x)^2-((1-x)-h)^2}{(1-x)^2((1-x)-h)^2h}\\ \, \\ =\lim_{h\rightarrow 0} \frac{(1-x)^2-[(1-x)^2-2h(1-x)+h^2]}{(1-x)^2[(1-x)^2-2h(1-x)+h^2]h}\\ \, \\ =\lim_{h\rightarrow 0} \frac{2h(1-x)-h^2}{(1-x)^2h[(1-x)^2-2h(1-x)+h^2]} \\ \, \\ =\lim_{h\rightarrow 0} \frac{\cancel{h}[2(1-x)-h]}{(1-x)^2\cancel{h}[(1-x)^2 -2h(1-x)+h^2]}\], plug in \(h=0\): \[ =\frac{2(1-x)}{(1-x)^4}\\ = \frac{2}{(1-x)^3}\]

OpenStudy (kirbykirby):

:)

OpenStudy (kirbykirby):

same answer as before so it's good

OpenStudy (anonymous):

Thanks!!

OpenStudy (kirbykirby):

yw